OpenStudy (wasiqss):

find among all possible straight lines through point(1,2), a line that forms a triangle of minimum area with the positive parts of the cartesian axes

OpenStudy (wasiqss):

i want minimum area

OpenStudy (eyust707):

what class is this for?

OpenStudy (wasiqss):

caclulus, application of maxima , minima

OpenStudy (wasiqss):

zarkon guide plz

OpenStudy (zarkon):

where are you stuck?

OpenStudy (wasiqss):

like i cant understand the whole question

OpenStudy (zarkon):

can you give me a general equation of a line that passes through the point (1,2)?

OpenStudy (wasiqss):

y-2=m(x-1)

OpenStudy (zarkon):

write is as y=m(x-1)+2 what the the x and y intercepts

OpenStudy (zarkon):

in terms of m

OpenStudy (wasiqss):

y intercept 2-m

OpenStudy (zarkon):

ok

OpenStudy (wasiqss):

i guess we have to find line with least gradient ,m i rite ?

OpenStudy (zarkon):

you need to minimize \[A=\frac{1}{2}base\times height\]

OpenStudy (zarkon):

the length of the base is the x intercept and the height of the triangle is the y-intercept

OpenStudy (wasiqss):

ok

OpenStudy (wasiqss):

w8 how base is x intercept and height y intercept

OpenStudy (zarkon):



OpenStudy (wasiqss):

ohhh yes now i understood, bt wut is next step

OpenStudy (zarkon):

you need the x-intercept

OpenStudy (wasiqss):

ok, we can do that by putting y=0 , the line equation